OpenStudy (anonymous):

What is the quotient in simplest form? State any restrictions on the variable.

OpenStudy (anonymous):

\[\frac{ z^2 - 4 }{z - 3} \div \frac{ z + 2}{ z^2 + z - 12 }\]

OpenStudy (anonymous):

\[\frac{ z^2 - 4 }{ z - 3 } \div \frac{ z+2 }{ z^2+z-12 }= \] \[\frac{ z^2 - 4 }{ z - 3 } * \frac{ z^2+z-12 }{ z+2 } = \] Factor the equations that need to be factored here. \[\frac{ (z+2)(z-2) }{ z - 3 } * \frac{ (z+4)(z-3) }{ z+2 } = \] And here you cancel out the stuff that can be canceled. \[(z-2) (z+4)\]